OpenStudy (anonymous):

Without graphing is the system independent, dependent, or inconsistent? -2x-y=9 3x-4y=-8 Could someone please explain

OpenStudy (experimentx):

is equation 2 multiple of equation 1 ?

OpenStudy (anonymous):

Not sure what you mean by a multiple

OpenStudy (anonymous):

Inconsistent equations are those which have no common solution: parallel lines. Dependent equations are 2 different ways of expressing the same line: infinite # of solutions. Independent equations have one ordered pair solution: they cross at one point.

OpenStudy (experimentx):

equation 1 -2x-y=9 equation 2 \( a \times (-2x-y)= a \times 9 \) if you have such type of system, then it is dependent.

OpenStudy (anonymous):

So, try to solve for y in terms of x in the first equation and substitute into the next equation and see where this takes you.

OpenStudy (anonymous):

I dont understand how to work the problems out to figure the answer, these are new to me

OpenStudy (experimentx):

if you don't have above type of systems then it is independent system.

OpenStudy (experimentx):

equation 1: \( −2x−y=9 \) equation 2: \( a×(−2x−y)=c\) where \( c≠a×9 \) if you have such type of system, then it is inconsistent..

OpenStudy (anonymous):

if i did it right i got y=-2x-9 and y=3x/4+2

OpenStudy (anonymous):

Good. Now, set both equation = to each other and solve for x.

OpenStudy (anonymous):

i ended up with x=-4 and y=-1

OpenStudy (anonymous):

Very good. You have a unique, one-point solution, so you have independent equations. And you're done!

OpenStudy (anonymous):

Picturing the system of the 2 equations in your mind's eye, since you're told not to graph, you can picture two lines that intersect at one point.

OpenStudy (anonymous):

if you have a problem that says no solution would that be considered inconsistent

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, thank you for helping me really appreciate it

OpenStudy (anonymous):

And if the system is of two variables, they are parallel lines on a flat plane in the Euclidean model, the one we think of automatically. The "real-world" model.

OpenStudy (anonymous):

You're quite welcome and thx for the medal recognition.

OpenStudy (anonymous):

your welcome